OpenStudy (ujjwal):

If xsin(a+y)=siny, then \[\frac{dy}{dx}=\] a)\[\frac{\sin^2 (a+y)}{\cos y}\] b)\[\frac{\sin a}{\cos^2 (a+y)}\] c)\[\frac{\sin^2 (a+y)}{\sin a}\] d)\[\frac{\sin a}{\sin^2 (a+y)}\]

OpenStudy (anonymous):

Take the derivative implicitly... Can you??

OpenStudy (ujjwal):

shouldn't it be equal to \[\frac{\sin (a+y)}{\cos y-x\cos (a+y)}\]

OpenStudy (anonymous):

Firstly you can separate y and x terms: \[x = \frac{\sin(y)}{\sin(a + y)}\] Now take the derivative implicitly: \[dx = \frac{\sin(a+y)(cosy) - \sin(y)(\cos(a + y)}{\sin^2(a+y)}dy\] \[\frac{dy}{dx} = \frac{\sin^2(a + y)}{\sin(a + y - y)} \implies \frac{dy}{dx} = \frac{\sin^2(a + y)}{\sin(a)}\]

OpenStudy (anonymous):

\[\sin(A - B) = sinAcosB - cosAsinB\] I have used this in the numerator..

OpenStudy (anonymous):

I mean numerator after solving will be equal to this..

OpenStudy (anonymous):

u could let \[x=\frac{\sin y}{\sin (a+y)} \] in what u got up there

OpenStudy (ujjwal):

yeah, i just did that.. That gives the result.. I don't know why i didn't do that earlier..